# A formal implementation of Behavior Trees to act in robotics

### v1

@@ 7.2 Discussion
Our approach to improve the trust one can put in BT completely relies on a on a well established formal language and framework (Fiacre and TTS), a formal verification toolbox (Tina to check LTL, CTL, properties) and Hippo, a runtime engine able to execute the formal model. We have seen in Section 2.2 that there are other approaches which transform BT in a formal framework, or harness a formal model around BT, but to our knowledge, none of them support the automatic translation of BT to an equivalent formal model and the execution the same formal model in place of the original BT. This is a very strong argument as the same formal model, which clarifies and specifies the formal semantics of BT, can be used to prove properties and also execute and show while running that it properly implements the expected operational semantic, and knowing that the proof made offline also holds online. As for taking into account the execution environment, it is taken into account in the perception provided by Condition and Action nodes.
Moreover, we have shown that the formal model of the various BT nodes (even the ones added for new experiments) can also be formally checked with our toolbox. Moreover, even if Fiacre and Tina do not provide per se a timed logic (e.g. TCTL or MITL), the toolbox provides some patterns Abid et al. (2014) which allow the verification of properties with explicit time (e.g. to prove that at most x units of time will elapse between these two states).
Moreover, Fiacre offers some features which could be valuable to be added to BT. Explicit time representation is the foremost feature which could come handy if added to the BT . Indeed, robot planning, acting and control is usually handling explicit time representation. Control loops have frequency, plans have duration and deadline, actions take some time to execute, etc. So it would be perfectly logical to add explicit time representation to BT. As a consequence, this would also have an impact on the tick semantic which would probably become more an execution token than a time elapsing counter.
Another domain where our approach and Fiacre could offer some valuable addition is with environment modelling. BT are intrinsically embedded operational models, they really exists to be executed in the real world (or in simulation). But now that we can consider model checking them, one needs to take into account (when possible) the outcomes of actions or conditions in the environment. Without any particular information, the model checker consider all possibilities (success, failure or running), but the real world often prove to be more “constrained”, and one could consider a more accurate and explicit model of the environment (e.g., with external asynchronous events, state variable values changes, etc).
@@
Last, our approach has been implemented in a software suite Ingrand (2024b) which has been applied to many BT examples (available here: https://redmine.laas.fr/projects/bt2fiacre/pages/index)
@@ 7.4 Conclusion
BT are more and more popular in robotics. To encourage their deployment and improve the trust one has in robotic applications using BT, we propose an approach and an automatic tool to transform any BT in a formal model with a formal semantics. The resulting models can then be used offline with model checking to prove some properties of the BT, but also linked to the real actions and perceptions of the robots and execute online on the robot. Of course, this can be deployed in BT applications outside of robotics, and also participates to define extensions to BT and to better formalize them.
@@ 5 Fallback27_btn13, Fallback27_btn13_done, done
Appendix C A ROS2 Nav2 BT and its BTF equivalent.
Listing 10: One of the ROS2 Nav2 Behavior Tree (Navigate To Pose With Replanning and Recovery) Macenski et al. (2024).
@@
Figure 18: The graphical representation of ROS2 Nav2 BT Listing 11p11.
Appendix D The Mars rover example from Biggar and Zamani (2020); Wang et al. (2024).
Listing 12: The BTF model of the Mars rover example from Biggar and Zamani (2020); Wang et al. (2024).

### v3

@@ 7.2 Discussion
Our approach to improve the trust one can put in BT completely relies on a on a well established formal language and framework (Fiacre and TTS), a formal verification toolbox (Tina to check LTL, CTL, properties, patterns Abid et al. (2014) [^19] Patterns allow the verification of properties with explicit time (e.g. to prove that at most x units of time will elapse between two states.).) and Hippo, a runtime engine able to execute the formal model. We have seen in Section 2.2 that there are other approaches which transform BT in a formal framework, or harness a formal model around BT, but to our knowledge, none of them support the automatic translation of BT to an equivalent formal model and the execution of the same formal model in place of the original BT. This is a very strong argument as the same formal model, which clarifies and specifies the formal semantics of BT, can be used to prove properties and also execute and show, while running, that it properly implements the expected operational semantic, moreover knowing that the proof made offline also holds online.
Despite the formal and proven tools deployed, our translation tool BT2Fiacre may still contain errors and bugs, but again, the fact that we can use theses tools to prove the resulting translated model against specifications done on the BT is reassuring. One could take every single BT node type, write its formal specification and prove that the Fiacre translation satisfies them with Tina (as we did for the ones added in the Nav2 experiment).
Furthermore, Fiacre offers some features which could be valuable to be added to BT. Explicit time representation is the foremost feature which could come handy if added to the BT. Indeed, robot planning, acting and control is usually handling explicit time representation. Control loops have frequency, plans have duration and deadline, actions take some time to execute, etc. So it would be perfectly logical to add explicit time representation to BT. As a consequence, this would also have an impact on the tick semantic which would probably become more an execution token than a time elapsing counter.
Another domain where our approach and Fiacre could offer some valuable addition is with environment modelling. BT are intrinsically embedded operational models, they really exist to be executed in the real world (or in simulation). But now that we can consider model checking them, one needs to take into account (when possible) the outcomes of actions or conditions in the environment. Without any particular information, the model checker consider all possibilities (success, failure or running), but the real world often prove to be more “constrained”, and one could consider a more accurate and explicit model of the environment (e.g., with external asynchronous events, state variable values changes, etc).
@@
Last, our approach has been implemented in a software suite Ingrand (2024b) which has been applied to many BT examples (available here: https://redmine.laas.fr/projects/bt2fiacre/pages/index). Moreover, all the examples presented here are available in the examples directory and we invite other formal approach to test their systems on these examples and report their results for comparison.
@@ 7.4 Conclusion
BT are more and more popular in robotics. To encourage their deployment and improve the trust one has in robotic applications using BT, we propose an approach and an automatic tool to transform any BT in a formal model with a formal semantics. The resulting models can then be used offline with model checking to prove some properties of the BT, but also linked to the real actions and perceptions of the robots and executed online on the robot. Of course, this can be deployed in BT applications outside of robotics, and also participates to define extensions to BT and to better formalize them.
@@ 5 Fallback27_btn13, Fallback27_btn13_done, done
Appendix C Fallback and Parallel nodes transformation in Fiacre
Figure 14: The Fiacre process modeling the Fallback node.
Figure 15: The Fiacre process modeling the Parallel node.
Appendix D The Drone BT.
Figure 16: The graphical representation of the BT Listing 4p4.
Appendix E A ROS2 Nav2 BT and its BTF equivalent.
Listing 10: One of the ROS2 Nav2 Behavior Tree (Navigate To Pose With Replanning and Recovery) Macenski et al. [2024].
@@ Appendix C Fallback and Parallel nodes transformation in Fiacre
Figure 17: The screen dump of the Nav2 BT (See Listing 11p11 executing.
Appendix F The Mars rover example from Biggar and Zamani [2020], Wang et al. [2024].
Listing 12: The BTF model of the Mars rover example from Biggar and Zamani [2020], Wang et al. [2024].